OpenStudy (anonymous):

Solve the equation: log(base 3)(x^2+1)=2?

OpenStudy (mimi_x3):

\(log_3(x^2+1)=2\)?

OpenStudy (anonymous):

yeah

OpenStudy (shubhamsrg):

well we indirectly have this : 3^2 = x^2 +1 will it be easier now ?

OpenStudy (anonymous):

yeah i got -2 root 2 and 2 root 2. Thanks!

OpenStudy (shubhamsrg):

glad to have helped ^_^